OpenStudy (anonymous):

Let R be the region bounded by the X- axis , the graph of y=Sqrt(X+1) and the line x=3.Find the area of the region R.

OpenStudy (campbell_st):

draw a picture 1st  the problem is now \[\int\limits_{-1}^{3} (x+1)^{1/2} dx\] the function is \[F(x) = 2/3 (x+1)^{3/2}\] last step F(3) - F(-1)